OpenStudy (anonymous):

use sqrt property to solve: (x-7)^2=8

OpenStudy (anonymous):

x = 7 + 2sqrt2 x = 7 - 2sqrt2

OpenStudy (anonymous):

(x-7)^2=8 \[x-7=\pm \sqrt{8}=\pm2\sqrt{2}\] x=\[7+2\sqrt2 , 7-2\sqrt2\]